1. Given: F(1,25). What is N?

2. Given: F(5,25). What is k?

3. Given: F(2,25). What is F(crit) for alpha = .01?
Answer #1

1)

F(k-1,n-k) = F(1,25)

So

k-1 =1 implies k = 2

n - k = 25

n = 27

2)

k-1 = 5

k = 6

3)

F(2,25) = 5.568 (From Critical value table)